OpenStudy (anonymous):

\[\sqrt[3]{x}^7 * \sqrt[3]{x}^2\]

jhonyy9 (jhonyy9):

cuberoot(x)9 =x3

jhonyy9 (jhonyy9):

\[\sqrt[3]{x ^{9}}=x ^{3}\]

jhonyy9 (jhonyy9):

ok ?

OpenStudy (mertsj):

\[x ^{\frac{7}{3}}x ^{\frac{2}{3}}\]

OpenStudy (mertsj):

\[x ^{\frac{9}{3}}=x ^{3}\]